### v2

# Beyond Sparsity: Quantum Block Encoding for Dense Matrices via Hierarchically Low Rank Compression

Abstract
While quantum algorithms for solving large scale systems of linear equations offer potential speedups, their application has largely been confined to sparse matrices. This work extends the scope of these algorithms to a broad class of structured dense matrices arise in potential theory, covariance modeling, and computational physics, namely, hierarchically block separable (HBS) matrices. We develop two distinct methods to make these systems amenable to quantum solvers. The first is a pre-processing approach that transforms the dense matrix into a larger but sparse format. The second is a direct block encoding scheme that recursively constructs the necessary oracles from the HBS structure. We provide a detailed complexity analysis and rigorous error bounds for both methods. Numerical experiments are presented to validate the effectiveness of our approaches.
I Introduction
Recent advances [aruteQuantumSupremacyUsing2019] in quantum computing have shown its potential for addressing problems that are difficult for classical computers. Especially, the prospect of exponential speedups has driven the exploration of quantum algorithms across a wide range of scientific and industrial areas. An important application of quantum computing is solving large scale linear systems. The first quantum algorithm for this purpose is the HHL algorithm [harrowQuantumAlgorithmSolving2009], which provides an exponential speedup for the quantum linear systems problem (QLSP). Subsequently, the quantum singular value transformation (QSVT) framework [gilyenQuantumSingularValue2019] was proposed, which can also be applied to solve linear systems. Readers are referred to [moralesQuantumLinearSystem2025] for a comprehensive review of these methods.
However, the performance of these algorithms relies on a key assumption: sparsity. The HHL algorithm, for example, depends on the ability to efficiently simulate the Hamiltonian evolution of the system matrix. The efficiency of this simulation, in turn, depends on the matrix’s sparsity [childsFirstQuantumSimulation2018]. Similarly, QSVT-based approaches require access to a block encoding of the system matrix. While constructing a block encoding for a sparse matrix is relatively straightforward, doing so for a dense matrix remains a challenge. This sparsity issue becomes severe because, under data access models similar to those used for block encodings, it is possible to construct classical randomized algorithms that match the performance of their quantum counterparts [gilyenImprovedQuantuminspiredAlgorithm2022]. As a result, better approaches for dense matrices are needed, but existing methods have their own drawbacks. For instance, the quantum algorithms studied in [wossnigQuantumLinearSystem2018] handle dense matrices with only a polynomial speedup instead of an exponential one.
The goal of this paper is to overcome the difficulty of constructing an efficient block encoding for dense matrices. We do so by exploiting the inherent special structures within certain important classes of dense matrices [nguyenBlockencodingDenseFullrank2022], including those arising in integral equations, covariance modeling, and computational physics. This structure often arises when matrix entries are defined by a kernel function $K(x,y)$, where interactions between well-separated clusters of points can be approximated by low rank operators. This principle gives rise to structured matrices like hierarchically off-diagonal low rank (HODLR) [laiFastDirectSolver2014] and hierarchically semiseparable (HSS) matrices [xiaFastAlgorithmsHierarchically2010]. HSS matrices employ a nested basis, where the basis at a given hierarchical level is constructed from those of its children. For more information of these hierarchical structures, see [hackbuschHierarchicalMatricesAlgorithms2015].
Motivated by the work of Ho and Greengard [hoFastDirectSolver2012], we proposes and analyzes two quantum approaches for solving dense linear systems based on hierarchical matrix representations. The first is an extended sparsification method that transforms the dense matrix into a larger, sparse system, making it amenable to efficient block encoding designed for sparse matrices. The second is a direct, recursive construction of the block encoding that follows the hierarchical structure. We provide a rigorous analysis for both strategies. For the sparsification approach, we show that the construction cost and resulting sparsity scale linearly with the matrix dimension. The condition number and approximation errors are also provided. For the recursive approach, we derive explicit bounds on the growth of the subnormalization factor and the accumulation of precision error.
The remainder of this paper is organized as follows. In Section 2, we review the basic formulations of quantum linear system algorithms and show the obstacles in constructing block encodings for dense matrices. Section 3 introduces HBS matrices and presents a modified algorithm adapted for the subsequential block encoding in quantum setting. In Section 4, we propose a sparsification based quantum approach, which uses an extended sparsification to combine the HBS matrix components into a single but slightly larger system. As an alternative, Section 5 describes a direct method for constructing the block encoding from the hierarchical matrix structure. Section 6 presents numerical results from experiments on integral equations to evaluate the performance of our methods based on the classical computers. Finally, Section 7 provides concluding remarks and discusses potential directions for future work.
II quantum linear system algorithms
To solve linear systems in the quantum setting, we first review the definition of the quantum linear system problem (QLSP) [childsQuantumAlgorithmSystems2017]. In the following presentation, if not explicitly mentioned, we always denote $\|\cdot\|$ the $2$-norm.
Definition II.1 (QLSP).
Consider an $N\times N$ Hermitian matrix $A$, which has condition number $\kappa$ and is normalized such that $\|A\|=1$. Let $\vec{b}$ be an $N$-dimensional vector and $\vec{x}:=A^{-1}\vec{b}$. The corresponding quantum states are defined as
[…]
Definition II.2 (Block Encoding).
An $(n+a)$-qubit unitary $U$ is an $(\alpha,a,\varepsilon)$-block-encoding of a $2^{n}\times 2^{n}$ matrix $A$ if it satisfies:
[…]
where $\alpha\geq\|A\|$ is the subnormalization factor, $a$ is the number of ancillary qubits, and $\varepsilon>0$ is a sufficiently small precision number.
The construction of block encodings can be achieved through several established methods, particularly for matrices with specific sparsity structures. We call a matrix $A$ is $s$-row-sparse ($s$-column-sparse) if each row (column) of $A$ has at most $s$ nonzero elements. The following lemma, adapted from Gilyén et al. [gilyenQuantumSingularValue2019] and summarized in Nguyen et al. [nguyenBlockencodingDenseFullrank2022], provides an example for oracle accessible sparse matrices.
Lemma II.1 (Block encoding [nguyenBlockencodingDenseFullrank2022]).
Let $A\in\mathbb{C}^{2^{n}\times 2^{n}}$ be a matrix that is $s_{r}$-row-sparse and $s_{c}$-column-sparse. Suppose we are given access to the following $(n+1)$-qubit oracles:
[…]
where $\tilde{a}_{ij}$ is the $b$-qubit fixed-point representation of $a_{ij}/\hat{a}$. Then, one can implement an $(\hat{a}\sqrt{s_{r}s_{c}},n+3,\varepsilon)$-block-encoding of $A$ with a single query to $\mathcal{O}_{r}$ and $\mathcal{O}_{c}$, two queries to $\mathcal{O}_{A}$, and $O(n+{\rm polylog}(\frac{\hat{a}\sqrt{s_{r}s_{c}}}{\varepsilon}))$ additional one and two qubit gates, using $O(b,{\rm polylog}(\frac{\hat{a}\sqrt{s_{r}s_{c}}}{\varepsilon}))$ ancilla qubits.
Lemma II.1 establishes that the efficiency of the block encoding, specifically the subnormalization factor $\alpha$, is determined by the geometric mean of the row and column sparsities. While this construction is highly efficient for sparse matrices, its limitations become apparent for dense matrices. For example, if $s_{r}=s_{c}=2^{n}$, the subnormalization factor $\alpha=\hat{a}\sqrt{s_{r}s_{c}}$ scales as $\hat{a}\cdot 2^{n}$. Since the complexity of algorithms using this block encoding typically scales polynomially with $\alpha$, this exponential growth in $n$ poses a significant drawback. We address this problem by proposing and analyzing two methods. The first uses an extended sparsification technique to convert the dense matrix into a sparse one, followed by an analysis of the cost to construct a block encoding for the resulting system. The second is a direct method for block encoding the hierarchical matrix structure itself, for which we analyze the complexity and the required subnormalization factor.
In the complexity analysis that follow, we account for the number of non-zero elements. This number provides a more realistic measure of resource requirements than simply using the matrix dimensions. The cost of implementing the oracles, such as $\mathcal{O}_{r},\mathcal{O}_{c}$ and $\mathcal{O}_{A}$ in Lemma II.1, is directly related to the gate count needed to access the value and location of each non-zero entry. In particular, [zhangCircuitComplexityQuantum2024] shows that the gate count for constructing these oracles is directly proportional to the number of non-zero elements in the matrix representation. Therefore, to implement our proposed methods effectively, we first require a suitable way to compress the dense matrix into a format with a manageable number of entries. The next section introduces the hierarchically block separable (HBS) format [hoFastSemidirectLeast2014] that serves as the foundation for our analysis.
III Hierarchically block separable matrices
Solving large scale linear systems is a classical problem in scientific computing. While standard iterative methods like Bi-CGSTAB and GMRES are often applied, their computational efficiency is matrix dependent [doi:10.1137/1.9781421407944], which can lead to poor performance for the dense matrices that arise in many physical problems. An effective strategy is to exploit the specific structure hidden behind these matrices. For instance, matrices from the discretization of integral equations are not arbitrarily dense but possess a sparse structure. This property stems from the smoothness of the underlying kernel $K(x,y)$ for well separated points $x$ and $y$ in 2D or 3D. Important kernels in potential theory and wave physics that exhibit this behavior include:
  1. 1.
    $\log|x-y|$, for 2D potential problems,
  2. 2.
    $H^{(1)}_{0}(\kappa|x-y|)$, for 2D wave scattering,
  3. 3.
    1/|xy|1/|x-y|, for 3D potential problems,
  4. 4.
    $e^{i\kappa|x-y|}/|x-y|$, for 3D wave scattering,
where $\kappa>0$ is the wavenumber, and $H^{(1)}_{0}$ is the first kind Hankel function of order zero.
The HBS matrix format provides an efficient way to represent this structure [hoFastDirectSolver2012, mindenRecursiveSkeletonizationFactorization2017]. To define this structure formally, consider a dense matrix $A\in\mathbb{C}^{N\times N}$. Let its row and column index sets be $I=\{1,\dots,N\}$ and $J=\{1,\dots,N\}$, respectively. We begin by partitioning the row index set $I$ into $p$ disjoint blocks, $\{I_{j}\}_{j=1}^{p}$. The $j$-th row block, $I_{j}$, contains $n_{j}=|I_{j}|$ indices and is defined as:
$I_{j}=\left\{p_{j-1}+1,p_{j-1}+2,\dots,p_{j}\right\}.$
where $p_{j}=\sum_{k=1}^{j}n_{k}$ represents the cumulative size of the first $j$ blocks. By convention, $p_{0}=0$. Similarly, the column index set $J$ is partitioned into $p$ blocks, $\{J_{k}\}_{k=1}^{p}$, where the $k$-th block, $J_{k}$, has a size of $n^{\prime}_{k}=|J_{k}|$. Using this partitioning, the linear system $Ax=b$ can be expressed in the block form:
$\sum_{j=1}^{p}A_{ij}x_{j}=b_{i},\quad\text{for }i=1,\dots,p,$
where $x_{j}\in\mathbb{C}^{n^{\prime}_{j}}$ and $b_{i}\in\mathbb{C}^{n_{i}}$ are the corresponding blocks of the vectors $x$ and $b$, and $A_{ij}\in\mathbb{C}^{n_{i}\times n^{\prime}_{j}}$ is the submatrix at the intersection of row block $I_{i}$ and column block $J_{j}$. Given an error tolerance $\varepsilon$, a matrix $A$ is defined as block separable if each of its off-diagonal blocks, $A_{ij}$ for $i\neq j$, admits a low-rank factorization of the form:
$A_{ij}\approx A_{ij,\varepsilon}=L_{i}S_{ij}R_{j}.$
Here, $L_{i}\in\mathbb{C}^{n_{i}\times k_{i}^{r}}$ and $R_{j}\in\mathbb{C}^{k_{j}^{c}\times n_{j}^{\prime}}$ are basis matrices, and $S_{ij}\in\mathbb{C}^{k_{i}^{r}\times k_{j}^{c}}$. The ranks $k_{i}^{r}$ and $k_{j}^{c}$ are assumed to be much more smaller than the block dimensions, i.e., $k_{i}^{r}\ll n_{i}$ and $k_{j}^{c}\ll n_{j}^{\prime}$.
This sparse representation can be constructed using the interpolative decomposition (ID) [halkoFindingStructureRandomness2011], which is defined as follows:
Definition III.1.
Let $A\in\mathbb{C}^{m\times n}$ be a matrix with a target rank $k<\min(m,n)$. A rank-$k$ ID of $A$ is a factorization of the form $A\approx A_{\varepsilon}=BP$ under a given error tolerance $\varepsilon$, where $B\in\mathbb{C}^{m\times k}$ is a skeleton matrix whose columns are a specific subset of $k$ columns from $A$. The matrix $P\in\mathbb{C}^{k\times n}$ is an interpolation matrix that contains the $k\times k$ identity matrix as a submatrix.
Efficient randomized algorithms can compute the ID with a typical complexity of $O(mn\log(k)+k^{2}n)$ [halkoFindingStructureRandomness2011]. ID is preferred over other low-rank methods like the singular value decomposition (SVD) for several reasons. Firstly, the skeleton matrix $B$ consists of actual columns from $A$, which preserves physical structure. Secondly, the interpolation matrix $P$ has a provably bounded norm, ensuring numerical stability. We use the ID algorithm to approximate the off-diagonal blocks $A_{ij}$ (for $i\neq j$) as $A_{ij}\approx L_{i}S_{ij}R_{j}$, where $S_{ij}\in\mathbb{C}^{k_{i}^{r}\times k_{j}^{c}}$ with ranks satisfying $k_{i}^{r}\ll n_{i}$ and $k_{j}^{c}\ll n_{j}$.
This recursive compression algorithm, illustrated in Figure 1, constructs the hierarchical factors ($D_{l},L_{l},R_{l}$) level by level. Let $S_{0}=A$ be the initial matrix. At each level $l$ of the hierarchy, the algorithm processes the current matrix $S_{l-1}$, which is partitioned into blocks based on the spatial tree. The blocks corresponding to near field, where the target $x$ and source $y$ are close, are typically dense, and these blocks are extracted and collectively form the diagonal factor $D_{l}$. The remaining matrix, which we denote as $B_{l}=S_{l-1}-D_{l}$, contains only the far field interactions. These far field blocks are presumed to be numerically low-rank. They are approximated using ID to identify a small set of skeleton indices and to construct basis matrices. To perform this efficiently, all far-field blocks within a given row partition share a common basis matrix, and similarly, all blocks within a column partition share a common column basis. The computational cost of these ID procedures can be significantly reduced by using proxy approximation, which replaces global computations with local ones involving a small set of sources [xingInterpolativeDecompositionProxy2020]; a more detailed description is provided in Appendix LABEL:ap:classic. These bases are then assembled into the block-diagonal factors $L_{l}$ and $R_{l}$. The far field part is thus approximated as $B_{l}\approx L_{l}S_{l}R_{l}$, where $S_{l}$ is the skeleton matrix. This new matrix $S_{l}$ is passed to the next level, $l+1$, for the same decomposition procedure. This process is repeated until a predefined recursion depth $\lambda$ is reached, and the final remaining skeleton matrix is set as $D_{\lambda+1}$.
Figure 1: This figure illustrates the algorithmic process of the recursive HBS factorization. At level 1, (a) the matrix $A$ is decomposed into its near-field blocks ($D_{1}$) and far-field remainder ($B_{1}$). (b) The far-field part is approximated as $B_{1}\approx L_{1}S_{1}R_{1}$ using ID. (c) The process is recursively applied to $S_{1}$, and the indices of the resulting far-field part $B_{2}$ are reordered for the subsequent level.
The recursive application of this compression is guided by a partition tree. A standard approach is to recursively bisect the row and column index sets, as shown in Figure 2. However, this index-based partitioning has a major limitation in higher dimensions. In 1D, nearby indices often correspond to physically close points, but this correspondence breaks down in 2D or 3D. Geometrically close points can have widely separated indices, which means their interaction blocks will appear off-diagonal. Such blocks are typically high-rank and cannot be compressed effectively [hoFastSemidirectLeast2014].
Figure 2: An index tree for the interval $[0,1]$. The interval is first partitioned into $0,\frac{1}{2}]$ and $\frac{1}{2},1]$, and the two resulting nodes are then partitioned further. Each node at each level corresponds to an interval.
To address this limitation, based on the HBS algorithm [hoFastDirectSolver2012], a more robust strategy uses the geometric coordinates of the underlying data points, which are typically available for matrices defined by a kernel function [coulierInverseFastMultipole2017]. Instead of partitioning indices based on their ordering, the hierarchy is constructed directly from these spatial coordinates, for instance, by employing a quadtree in 2D or an octree in 3D. For matrices that are not strictly generated by a kernel function, precise coordinates are not strictly required for this partitioning; approximate coordinates are sufficient. For example, if the matrix represents a kernel-based system subject to a small perturbation, the coordinates associated with the original kernel function can still be effectively used. This hierarchical structure ensures that matrix blocks corresponding to well-separated spatial clusters are low rank and compressible, whereas blocks corresponding to adjacent points are generally treated as dense diagonal blocks at a given level. The procedure is summarized in Algorithm III and illustrated in Figure 1.
Algorithm 1 Construction of HBS Factors via Spatial Partitioning
0: Matrix $A\in\mathbb{C}^{N\times N}$; A spatial tree $\mathcal{T}$ partitioning the indices $\{1,\dots,N\}$; Recursion depth $\lambda$; Compression tolerance $\varepsilon$.
0: Diagonal blocks $D=\{D_{1},\dots,D_{\lambda},D_{\lambda+1}\}$; Row basis matrices $L=\{L_{1},\dots,L_{\lambda}\}$; Column basis matrices $R=\{R_{1},\dots,R_{\lambda}\}$.
[…]
4:for $k=1$ to $\lambda$ do
5:  Partition $I_{active}$ and $J_{active}$ into blocks $\{\mathcal{I}_{p}\}$ and $\{\mathcal{J}_{q}\}$ based on level $k$ of the tree $\mathcal{T}$.
6:  // Extract near-field blocks to form $D_{k}$
7:  $D_{k}\leftarrow$ Submatrix of $S$ formed by all near-field blocks at level $k$.
8:  $S\leftarrow S-D_{k}$ // Remove the diagonal elements
9:  // Compress blocks to find skeletons and bases
10:  $I_{\rm{skel}}\leftarrow\emptyset$, $J_{\rm{skel}}\leftarrow\emptyset$, $U_{\rm{list}}\leftarrow\{\}$, $V_{\rm{list}}\leftarrow\{\}$
11:  for each row block $\mathcal{I}_{p}$ do
12:   Apply ID to the submatrix $S(\mathcal{I}_{p},\mathcal{J}^{(p)}_{\rm{far}})$ with tolerance $\varepsilon$.
13:   Let $\mathcal{I}^{\prime}_{p}\subseteq\mathcal{I}_{p}$ be the skeleton row indices and $U_{p}$ be the corresponding basis matrix.
14:   $I_{\rm{skel}}\leftarrow I_{\rm{skel}}\cup\mathcal{I}^{\prime}_{p}$
15:   Append $U_{p}$ to $U_{\rm list}$.
16:  end for
17:  for each column block $\mathcal{J}_{q}$ do
18:   Apply ID to the submatrix $S(\mathcal{I}^{(q)}_{\rm{far}},\mathcal{J}_{q})$ with tolerance $\varepsilon$.
19:   Let $\mathcal{J}^{\prime}_{q}\subseteq\mathcal{J}_{q}$ be the skeleton column indices and $V_{q}$ be the corresponding basis matrix.
20:   $J_{\rm{skel}}\leftarrow J_{\rm{skel}}\cup\mathcal{J}^{\prime}_{q}$
21:   Append $V_{q}$ to $V_{\rm{list}}$.
22:  end for
23:  // Assemble factors and update for the next level
24:  $L_{k}\leftarrow\text{blkdiag}(U_{p}\in U_{\rm{list}})$ // Assemble row bases into a block-diagonal matrix
25:  $R_{k}\leftarrow\text{blkdiag}(V_{q}\in V_{\rm{list}})$ // Assemble column bases into a block-diagonal matrix
26:  $S\leftarrow S(I_{\rm{skel}},J_{\rm{skel}})$ // Update S to be the skeleton matrix
27:  $I_{\rm{active}}\leftarrow I_{\rm{skel}}$, $J_{\rm{active}}\leftarrow J_{\rm{skel}}$
28:end for
[…]
This process yields the telescoping representation
$A\approx D_{1}+L_{1}[D_{2}+L_{2}(\cdots D_{\lambda}+L_{\lambda}D_{\lambda+1}R_{\lambda}\cdots)R_{2}]R_{1},$ (1)
where $\lambda$ is the maximum depth of the hierarchy. The recursion is typically terminated when the matrix blocks at the final level reach a predefined size $M$. It determines the depth as $\lambda=O(\log N)$.
To analyze the complexity and sparsity of the algorithm, denote each level $l=1,\dots,\lambda$, where $l=1$ corresponds to the finest partition. Let $p_{l}$ be the number of blocks along one dimension, $n_{l}$ be the uncompressed block size, and $k_{l}$ be the compressed block size (i.e., the rank). For simplicity, we assume these parameters are uniform for all blocks at a given level. The spatial dimension is denoted by $d$. Our analysis is based on the following assumptions:
  1. (i)
    The total matrix dimension is $N=p_{1}n_{1}$. We assume the finest-level blocks contain a constant number of points $M$, so $n_{1}=O(1)$, which implies $p_{1}=O(N)$. For simplicity, we set $p_{1}=2^{d\lambda}$.
  2. (ii)
    At each level of the hierarchy, a parent block is divided into $2^{d}$ children. This means the number of blocks is divided by $2^{d}$ at each coarser level, so $p_{l}=p_{1}/2^{d(l-1)}$. Since $p_{\lambda}=O(1)$, the total number of levels is $\lambda=\lfloor\log_{2}(N/M)/d\rfloor+1=O(\log N)$.
  3. (iii)
    The size of the matrix at level $l$ is determined by the number of skeleton points from level l1l-1. This gives the relation $p_{l}n_{l}=p_{l-1}k_{l-1}$, so $n_{l}=2^{d}k_{l-1}$.
  4. (iv)
    The ranks of the off-diagonal blocks corresponding to well-separated points are bounded by a constant $r$ that depends on the desired precision but not on the matrix size $N$. We therefore assume $k_{l}=r$ for all levels $l$.
Assumption (iv) is justified for matrices generated by smooth kernel functions $K(x,y)$. The Taylor expansion of such a kernel shows that for well-separated domains, it can be approximated by a finite sum of separable functions:
$K(x,y)$ $\approx\sum\limits_{|{\bm{\beta}}|=0}^{r-1}\frac{(y-y_{0})^{\bm{\beta}}}{{\bm{\beta}}!}D^{\bm{\beta}}_{2}K(x,y_{0})$ (2)
$\approx\sum\limits_{|{\bm{\alpha}}|=0}^{r-1}\sum\limits_{|{\bm{\beta}}|=0}^{r-1}\frac{(x-x_{0})^{\bm{\alpha}}}{{\bm{\alpha}}!}\frac{(y-y_{0})^{\bm{\beta}}}{{\bm{\beta}}!}D_{1}^{\bm{\alpha}}D^{\bm{\beta}}_{2}K(x_{0},y_{0}).$
Truncating this expansion yields a low-rank approximation where the rank $r$ is determined by the required accuracy. This is the same underlying principle that enables the efficiency of algorithms like the fast multipole method [bebendorfHierarchicalMatricesMeans2008].
The classical procedure for constructing the decomposition of matrix $A$ leads to the following result regarding its computational cost and sparsity.
Theorem III.1.
Let $A\in\mathbb{C}^{N\times N}$ be a matrix that has a hierarchical representation of depth $\lambda$. If this representation is built using ID and a strong rank-revealing QR (RRQR) algorithm with parameter $f$, then the construction has a computational cost of
$T\sim 2^{d+1}(mr\log r+r^{3})(2^{d(\lambda+1)}-2^{d})/(2^{d}-1)\sim O(N),$ (3)
where $r$ is the maximum rank of the decompositions and $m$ is the proxy set size. The entries of the resulting matrices are bounded by a constant $c$. The matrices are sparse, containing a total of
$s=3^{d}n_{1}N+2^{d}[(2^{d+1}r^{2}-2r^{2}+4r)(2^{d\lambda}-1)+4^{d}(6^{d}-3^{d})r^{2}(2^{d\lambda}-1)]/(2^{d}-1)\sim O(N)$ (4)
non-zero elements.
The proof of Theorem III.1 can be found in Appendix LABEL:ap:classic. This theorem shows that the classical preprocessing cost and the number of non-zero elements in the resulting matrices both scale linearly with the matrix dimension $N$. Furthermore, the entries of these matrices are bounded. These properties are important considerations for quantum algorithms, and we will apply the results of Theorem III.1 in the analysis that follows.
IV A Sparsification Based Quantum Approach
The decomposition algorithm from Section III provides an effective method for creating a sparse representation of a matrix. However, this process yields a set of matrices, whereas a block encoding requires a single matrix. To address this, we use an extended sparsification technique to combine these matrices into a single but larger sparse matrix.
The HBS decomposition yields factors $L,D,$ and $R$ that are either sparse or much smaller than the original matrix. This property holds at each level of the hierarchy, allowing us to approximate the original matrix $A$ with controlled accuracy. We begin by considering a single level of compression. In this case, the matrix $A$ is approximated as $A\approx D+LSR$. The linear system $Ax=b$ can then be solved by replacing $A$ with this approximation:
$(D+LSR)x=b.$
To solve this system, we introduce the auxiliary variables $z=Rx$ and $y=Sz$, which allows us to write the system in the form
$\begin{pmatrix}D&L&\\ R&&-I\\ &-I&S\\ \end{pmatrix}\begin{pmatrix}x\\ y\\ z\end{pmatrix}=\begin{pmatrix}b\\ 0\\ 0\end{pmatrix}$.
Applying this procedure recursively to the telescoping representation in Equation (1) yields the following extended sparse system:
$\begin{pmatrix}D_{1}&L_{1}&&&&&\\ R_{1}&&-I&&&&\\ &-I&D_{2}&L_{2}&&\\ &&R_{2}&\ddots&\ddots&&\\ &&&\ddots&D_{\lambda}&L_{\lambda}&\\ &&&&R_{\lambda}&&-I\\ &&&&&-I&D_{\lambda+1}\\ \end{pmatrix}\begin{pmatrix}x\\ y_{1}\\ z_{1}\\ \vdots\\ \vdots\\ y_{\lambda}\\ z_{\lambda}\\ \end{pmatrix}=\begin{pmatrix}b\\ 0\\ 0\\ \vdots\\ \vdots\\ 0\\ 0\\ \end{pmatrix}$. (5)
We denote the block matrix in this system by $A_{\text{sp}}$ and refer to it as the extended sparsification of $A$. This structure has been employed in [hoFastDirectSolver2012] for the construction of a fast direct solver for boundary integral equations.
To be used in a quantum algorithm, we must analyze the properties of $A_{\text{sp}}$. As shown in Theorem III.1, the total number of non-zero elements in $A_{\text{sp}}$ scales as $O(N)$. We now consider its row and column sparsity. The non-zero entries in any given row of $A_{\text{sp}}$ come from at most three types of block matrices: a diagonal block $D_{l}$, the interpolation matrices $L_{l}$ and $R_{l-1}$, and an identity matrix $I$. By construction, the interpolation matrices at level $l$ are sparse, with at most $k_{l}$ or $k_{l}+1$ non-zeros per row or column. The diagonal blocks $D_{l}$ contain near-neighbor interactions. At the finest level ($l=1$), a block interacts with its $3^{d}-1$ neighbors, so each row in $D_{1}$ has at most $3^{d}n_{1}$ non-zeros. For coarser levels ($l>1$), a block interacts with its $(6^{d}-3^{d})$ neighbors, resulting in at most $(6^{d}-3^{d})n_{l}$ non-zeros per row in $D_{l}$. Combining these bounds, the maximum number of non-zeros in any row of $A_{\text{sp}}$ is
$s_{r}=\max\{3^{d}n_{1}+k_{1},n_{1}-k_{1}+1,(6^{d}-3^{d})k_{1}+k_{2}+1,n_{2}-k_{2}+1,\cdots,n_{\lambda}-k_{\lambda}+1,(6^{d}-3^{d})k_{\lambda}+1\}.$ (6)
Since $n_{l}$ and $k_{l}$ are constants for all levels, the row sparsity $s_{r}$ is bounded by a constant. A similar argument holds for the column sparsity $s_{c}$.
Finally, we consider the dimension of the matrix $A_{\text{sp}}$. The total number of rows, $N_{sp}$, is the sum of the dimensions of all the subblocks, which also scales linearly with $N$:
$N_{\text{sp}}=\sum\limits_{l=1}^{\lambda}2p_{l}k_{l}+p_{1}n_{1}\leq 2r(2^{d(\lambda+1)}-2^{d})+N\sim O(N).$ (7)
With these properties established, we can determine the cost of creating a block encoding for $A_{\text{sp}}$. Let $c_{\text{sp}}$ be the upper bound on the magnitude of elements in the original matrix $A$, and let $f$ be the parameter defined in Theorem LABEL:apth:RRQR for the strong RRQR algorithm [guEfficientAlgorithmsComputing1996]. Based on the conditions of Theorem II.1, we arrive at the following theorem:
Theorem IV.1.
Let $A$ be a matrix represented hierarchically to depth $\lambda$ with the extended sparsification $A_{\rm{sp}}=(a_{ij})_{N_{\rm{sp}}\times N_{\rm{sp}}}$ that is $s_{r}$-row-sparse and $s_{c}$-column-sparse. Then, one can implement an $(c_{\rm{sp}}\sqrt{s_{r}s_{c}},O(n),\varepsilon)$ block encoding of $A_{\rm{sp}}$ with a single query to $\mathcal{O}_{r}$ and $\mathcal{O}_{c}$, two queries to $\mathcal{O}_{A}$, and $O(n+{\rm polylog}(\frac{c_{\rm sp}\sqrt{s_{r}s_{c}}}{\varepsilon}))$ additional one and two qubit gates, using $O(b,{\rm polylog}(\frac{c_{\rm sp}\sqrt{s_{r}s_{c}}}{\varepsilon}))$ ancilla qubits. The oracles $O_{r},O_{c},O_{A}$ provide access to $O(N)$ nonzero matrix entries, where $N$ is the dimension of $A$.
While this sparsification method allows us to handle the dense matrix with algorithms designed for sparse systems, it also introduces three challenges. First, the solution vector of the extended system contains not only the desired solution $x$ but also the auxiliary variables $y_{l}$ and $z_{l}$. These additional components reduce the probability of measuring $x$, which increases the overall runtime of the algorithm. Second, the complexity of quantum linear system solvers is sensitive to the condition number of the matrix. The relationship between the condition number of the extended matrix $A_{\text{sp}}$ and the original matrix $A$ is not obvious, and the transformation could potentially make it much larger. Finally, the HBS representation is an approximation of the original dense matrix, and the impact of this initial error on the accuracy of the final quantum solution needs to be analyzed. Theorem IV.2 provides quantitative bounds for each of these issues.
Theorem IV.2.
Let $A\in\mathbb{C}^{N\times N}$ be a matrix approximated by a hierarchical matrix $A_{\varepsilon}$ with relative error $\|A-A_{\varepsilon}\|/\|A\|\leq\varepsilon$. Let $A_{\rm{sp}}$ be the extended sparse matrix derived from $A_{\varepsilon}$ with row sparsity $s_{r}$, column sparsity $s_{c}$, and entries bounded by $|(A_{\rm{sp}})_{ij}|\leq c_{\rm{sp}}$. The following bounds hold:
  1. (1)
    By applying a scaling with parameter $0<t<1$ in the hierarchical decomposition, the subnormalization factor of the resulting block encoding for the new sparse matrix $A^{\prime}_{\rm{sp}}$, which is obtained by replacing $L_{l}$ with $t^{-1}L_{l}$ and $R_{l}$ with $tR_{l}$ in $A_{\rm{sp}}$, is bounded by $\alpha(A^{\prime}_{\rm{sp}})\leq t^{-1}c_{\rm{sp}}\sqrt{s_{r}s_{c}}$. The squared norm of the auxiliary variables $(y^{\prime},z^{\prime})$ in the scaled solution $x^{\prime}=(x^{T},(y^{\prime})^{T},(z^{\prime})^{T})^{T}$ is constrained by the following inequality:
    $\sum_{l=1}^{\lambda}(\|y^{\prime}_{l}\|^{2}+\|z^{\prime}_{l}\|^{2})\leq t^{2}\sum_{l=1}^{\lambda}(\|y_{l}\|^{2}+\|z_{l}\|^{2}).$
  2. (2)
    For any regularization parameter $\alpha>0$, the condition number of the Tikhonov regularized normal matrix $A_{\rm{sp}}^{\dagger}A_{\rm{sp}}+\alpha I$ is bounded by:
    $\kappa(A_{\rm{sp}}^{\dagger}A_{\rm{sp}}+\alpha I)\leq\frac{c_{\rm{sp}}^{2}s_{r}s_{c}+\alpha}{\alpha}.$ (8)
  3. (3)
    Let $x=A^{-1}b$ and $x_{\varepsilon}=A_{\varepsilon}^{-1}b$. If the approximation is sufficiently accurate such that $\varepsilon\kappa(A)<1$, the relative error in the solution is bounded by:
    $\frac{\|x-x_{\varepsilon}\|}{\|x\|}\leq\frac{\varepsilon\kappa(A)}{1-\varepsilon\kappa(A)}.$ (9)
Details of Theorem IV.2 is proved in Appendix LABEL:ap:quantum. Theorem IV.2 shows that the costs associated with this sparsification method are controllable. The condition number can be bounded with regularization, and the approximation error is tied to the properties of the original system. These benefits come at a quantifiable cost in the success probability of the measurement.
V A Direct Block Encoding Algorithm for Hierarchical Matrices
An alternative to the extended sparsification method is to construct a block encoding of the matrix directly from its HBS approximation. Instead of forming an extended sparse system to solve a linear equation, this approach embeds the matrix itself into a unitary operator, as defined in Definition II.2. This method requires oracles that access only $O(N)$ nonzero elements, in contrast to the $O(N^{2})$ elements required by the approach in [nguyenBlockencodingDenseFullrank2022], and it works for a broader class of kernel functions. The construction follows the hierarchical structure of the matrix shown in Equation (1), proceeding from the lowest level upwards to build the encoding for the full matrix.
The construction relies on two main quantum primitives: the linear combination of unitaries (LCU) and quantum matrix multiplication (QMM). See Figures LABEL:fig:lcu_circuit and LABEL:fig:qmm_circuit for their quantum circuits, respectively. We assume that block encodings for the component matrices $D_{l}$, $L_{l}$, and $R_{l}$ are given, denoted as $U^{D}_{l}$, $U^{L}_{l}$, and $U_{l}^{R}$, respectively. Each encoding uses $a$ ancilla qubits and has an error of $\varepsilon$. The subnormalization factors for the diagonal blocks $D_{l}$ are denoted $\alpha_{l}^{(D)}$, and for the interpolation matrices $L_{l},R_{l}$ they are $\alpha_{l}^{(L)}$. Based on Lemma II.1 for block encoding sparse matrices, and recalling that the elements of $D_{l}$ are drawn from $A$, we have
$\alpha_{l}^{(D)}\leq(6^{d}-3^{d})n_{l}c_{\text{sp}},\ l>1,$ (10)
$\alpha_{1}^{(D)}\leq 3^{d}n_{1}c_{\text{sp}}.$
For the interpolation matrices, the subnormalization factor is bounded by
$\alpha_{l}^{(R)}=\alpha_{l}^{(L)}\leq f\sqrt{k_{l}(n_{l}-k_{l}+1)}.$ (11)
Since $n_{l}$ and $k_{l}$ are bounded by constants, these subnormalization factors are also bounded. To combine these block encoded matrices using LCU, we first need the concept of a state preparation pair.
Definition V.1 (State preparation pair).
[…]
The LCU technique allows for the construction of a linear combination of unitary operators, which in turn implements the addition of the matrices they block encode.
Lemma V.1 (LCU [gilyenQuantumSingularValue2019]).
[…]